OpenStudy (anonymous):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b plus one equals thirty-one Step 1: fifteen times b plus one minus one equals thirty-one minus one Step 2: fifteen times b equals thirty Step 3: fifteen times b times fifteen equals thirty times fifteen Step 4: b equals four hundred fifty

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

15b + 1 = 31 -- subtract 1 from both sides 15b + 1 - 1 = 31 - 1 -- combine like terms 15b = 30 -- He went wrong here. He should have divided both sides by 15 instead of multiplying both sides by 15. 15b/15 = 30/15 b = 2

OpenStudy (texaschic101):

He started messing up at step 3